OpenStudy (austinmcgowan196):

what is the center and radius of the circle: x^2 + y^2 - 6x + 4y = 12 medal could be yours

OpenStudy (mww):

First group the x's and y's separately then complete the square for both x and y

OpenStudy (mathmate):

The standard equation of a circle is \((x-x0)^2+(y-y0)^2=r^2\) .................(1) where (x0,y0) is the centre of the circle, and r is the radius. The equation of the given circle can be normalized as follows: \((x-3)^2+(y+2)^2=3^2+2^2=13\) Compare with equation (1) to find the centre (x0,y0), and the radius r, if necessary.

OpenStudy (mathmate):

addendum: Read following for how to normalize the equation of a circle, and a lot more information to help you. http://www.regentsprep.org/regents/math/algtrig/atc1/circlelesson.htm